OpenStudy (anonymous):

you have a 5 gallon and 3 gallon jug. Is it possible to make EXACTLY 4 gallons without using anything besides the 2 jugs and water. there are no markings on the jugs by which you can measure amounts less than 3 or 5 gallons.

OpenStudy (anonymous):

fill the 5 gall half way =2.5 and the 3 gall half way =1.5 2.5+1.5=4

jimthompson5910 (jim_thompson5910):

step 1) fill up the 5 gallon jug completely step 2) dump the contents of the 5 gallon jug into the 3 gallon jug. You will have 5-3 = 2 gallons left in the 5 gallon jug step 3) toss out the contents of the 3 gallon jug. step 4) pour the water from the 5 gallon jug (the 2 gallons left over) into the 3 gallon jug. There is 3-2 = 1 gallon left to fill in this 3 gallon jug step 5) fill up the 5 gallon jug completely step 6) pour the water from the 5 gallon jug into the 3 gallon jug to fill up the 3 gallon jug completely. You need 1 gallon of water to go from 2 gallons to 3 gallons. Taking 1 gallon from the 5 gallons gives you 5-1 = 4 gallons